A sample of CO2 occupies 225 ml at 373 K. At what temperature (in K) would it occupy 300 ml if the pressure did not change?



Grade:

1 Answers

Anurag Kishore
14 years ago

Hi, Cane

use the ideal gas equation

V / T = constant

225 / 373 = 300 / x

=>  x = 497.33 ml

 

Thanks

Anurag